# Corner modes in Non-Hermitian long-range model

### v1

@@ Abstract
We consider non-Hermitian (NH) analog of a second-order topological insulator, protected by chiral symmetry, in the presence of second-nearest neighbor hopping elements to theoretically investigate the interplay between long-range and topological order away from Hermiticity. In addition to the four zero-energy corner modes present in the first nearest neighbor hopping model, we uncover that the second nearest neighbor hopping introduces another topological phase with sixteen zero-energy corner modes. Importantly, the NH effects are manifested in altering the Hermitian phase boundaries for both the models. While comparing the complex energy spectrum under open boundary conditions, and bi-orthogonalized quadrupolar winding number (QWN) in real space, we resolve the apparent anomaly in the bulk boundary correspondence of the NH system as compared to the Hermitian counterpart by incorporating the effect of non-Bloch form of momentum into the mass term. The above invariant is also capable of capturing the phase boundaries between the two different topological phases where the degeneracy of the corner modes is evident, as exclusively observed for the second nearest neighbor model.
@@ I Introduction
The systems with topological band properties are identified with gapless boundary modes that are characterized by symmetry-protected topological invariants. This is known as bulk boundary correspondence (BBC) Hasan and Kane (2010); Sato and Ando (2017). The conventional BBC, being an integral part of the first-order ($n=1$) topological phase Hasan and Kane (2010); Haldane (1988); Kane and Mele (2005); Bernevig et al. (2006), is generalized for higher-order ($n>1$) topological phases in $d\geq 2$ dimensions where there exist $n_{c}=(d-n)$-dimensional boundary modes Benalcazar et al. (2017a, b); Song et al. (2017); Langbehn et al. (2017); Schindler et al. (2018); Franca et al. (2018); Wang et al. (2019); Călugăru et al. (2019); Szumniak et al. (2020); Ni et al. (2020); Xie et al. (2021); Saha et al. (2022); Zhu (2018); Liu et al. (2018); Yan et al. (2018); Wang et al. (2018a); Zhang et al. (2019a, b); Volpez et al. (2019); Yan (2019a); Ghorashi et al. (2019); Wu et al. (2020); Laubscher et al. (2020); Roy (2020); Zhang et al. (2020); Kheirkhah et al. (2021); Yan (2019b); Ahn and Yang (2020); Luo et al. (2021); Wang et al. (2018b); Ghosh et al. (2021); Roy and Juričić (2021); Li et al. (2021); Ghosh et al. (2022a, b). For example, the second-order topological insulator (SOTI) hosts zero-dimensional (0D) localized corner modes at zero-energy while this phase is characterized by nested polarization or quadrupolar moment. Very recently, it has been reported that the number of boundary modes in a topological phase can be tuned by considering long-range hopping terms DeGottardi et al. (2013); Niu et al. (2012); Xie et al. (2019); Dias and Marques (2022) as well as by implementing periodic Floquet drive Thakurathi et al. (2013); Ghosh et al. (2022a). Once the extended model continues to preserve the chiral symmetry (CS), one can characterize the new topological phase by winding numbers in odd spatial dimension  Schnyder et al. (2008); Kitaev (2009); Ryu et al. (2010). The number of degenerate zero-energy states at each boundary increases according to the enhancement of the range of the hopping amplitudes as indicated by the winding number ensuring the BBC Ryu and Hatsugai (2002); Teo and Kane (2010). It is noteworthy that one-dimensional winding number for the first-order topological systems becomes passive in case of even-dimensional generalizations Nakahara (2018). In contrast, the higher-order topological (HOT) phase in even spatial dimension can be characterized by an appropriately defined winding number preserving CS as a constraint Benalcazar and Cerjan (2022).
@@
In recent years, thanks to the practical realization of higher-order topological phases in meta-materials Parappurath et al. (2020); Yang et al. (2019); Malzard et al. (2015); Regensburger et al. (2012) where energy conservation no longer holds El-Ganainy et al. (2018); Denner et al. (2021) and the domain of topological quantum matter can be extended to the non-Hermitian (NH) systems. The coupling to the environment  Bergholtz and Budich (2019); Yang et al. (2021); San-Jose et al. (2016), disorder/interaction-mediated quasiparticles with finite lifetime Kozii and Fu ; Yoshida et al. (2018); Shen and Fu (2018) can effectively induce a complex self energy that is modeled by an NH effective Hamiltonian  Musslimani et al. (2008); Makris et al. (2008); Yao et al. (2018); Kawabata et al. (2019); Bergholtz et al. (2021); Sone et al. (2020); Denner et al. (2021); Ghosh and Black-Schaffer . Interestingly, the non-Bloch nature of the wavefunction for the NH systems renormalizes the topological mass term, thus enriching the BBC such that topological phase transitions perceived with open-boundary conditions can be explained by an appropriate bulk invariant  Yao and Wang (2018); Kunst et al. (2018); Liu et al. (2019); Helbig et al. (2020); Borgnia et al. (2020); Koch and Budich (2020); Zirnstein et al. (2021); Takane (2021). The NH topological systems showcase various intriguing features such as the skin effect where the bulk states accumulate at the boundary  Yao and Wang (2018); Kunst et al. (2018); Helbig et al. (2020); Kawabata et al. (2020), exceptional points where eigenstates, corresponding to the degenerate bands, coalesce Bender (2007); Heiss (2012).
@@
Going beyond the scope of short-range hopping, long-range hopping in Hermitian systems is found to mediate versatile topological phases where the number of zero-energy modes increases Pérez-González et al. (2019); Dias and Marques (2022); An et al. (2018); Hsu and Chen (2020). In this context, the interplay between long-range hopping elements and non-Hermiticity is still in its infancy as far as HOT systems are concerned  Liu et al. (2019); Ghosh and Nag (2022); Liu et al. (2021, 2023); Mondal and Nag . Therefore, considering a two-dimensional (2D) long-range NH SOTI, we, therefore, examine whether non-Hermiticity induces exceptional SOTI phases, otherwise absent in the Hermitian case, and address the following interesting questions that have not been explored so far in literature, to the best of our knowledge: How does the BBC change in the above NH phases? Can we characterize these emerging exceptional topological phases by bi-orthogonalized non-Bloch winding number?
@@
We consider a CS-preserved generic model, hosting SOTI phases in the presence of non-Hermiticity and long-range hopping terms. The first [second] nearest neighbor (NN) Hermitian model Hamiltonian can host four [four and sixteen] zero-energy corner modes while exceptional points, caused by the NH effects, reshape topological phase boundaries as compared to the Hermitian case. As a result, we find that BBC is not only different from their Hermitian counterparts but also non-trivially modified once long-range terms are included. The phase boundaries between four [one] and one [none] zero-energy modes per corner in the second [first] NN model are revised following the dressed mass term due to the non-Bloch nature of wavefunction for NH Hamiltonian. We compute the quadrupolar winding number (QWN) appropriately in real space by exploiting the CS as well as implementing the bi-orthogonalization and non-Bloch nature to lay out the phase diagram which is in accordance with the complex spectrum under open boundary conditions (OBC).
@@
The remainder of the article is organized in the following way. We discuss the details of the tight-binding Hamiltonian in Sec. II, where both the NH short-range and relatively long-range models are demonstrated. Sec. III is devoted to the main results of this article. In particular, we discuss the result associated with the first and the second NN neighbor hopping models in Secs. III.1 and III.2, respectively. Next, we illustrate the exceptional phase diagram in Sec. III.3 by examining the QWN. We finally summarize and conclude in Sec. IV.
@@ II Model Hamiltonian
where, $\Gamma_{1}=\sigma_{x}s_{z}$, $\Gamma_{2}=\sigma_{y}s_{0}$, $\Gamma_{3}=\sigma_{z}s_{0}$, $\Gamma_{4}=\sigma_{x}s_{x}$. We consider strengths of first (second) NN hopping, spin-orbit coupling, and $C_{4}$ symmetry breaking mass terms as $\lambda^{h}_{1}$, $\lambda^{s}_{1}$, and $\lambda^{f}_{1}$, ($\lambda^{h}_{2}$, $\lambda^{s}_{2}$, and $\lambda^{f}_{2}$) respectively. Here, $m_{0}$ is the staggered mass term. In what follows, we refer to the case of $\lambda_{1}\neq 0$ and $\lambda_{2}=0$ ($\lambda_{1,2}\neq 0$) as first (second) NN model. The SOTI phase with four zero-energy corner modes arises for $-2\lambda^{h}_{1}<m_{0}<2\lambda^{h}_{1}$ when $(\lambda^{f}_{1},\lambda^{f}_{2}=0)$. The SOTI phases with four and sixteen zero-energy corner modes exist for $-2\lambda^{h}_{2}<m_{0}<0$ and $0<m_{0}<2\lambda^{h}_{1}+2\lambda^{h}_{2}$, respectively when $\lambda^{f}_{1,2}\neq 0$. Note that, this model also hosts first order topological insulator phase for $-2\lambda^{h}_{1}<m_{0}<2\lambda^{h}_{1}$ and $-2\lambda^{h}_{2}<m_{0}<2\lambda^{h}_{1}+2\lambda^{h}_{2}$, respectively, when $(\lambda^{f}_{1},\lambda^{f}_{2})=(\neq 0,0)$ and (0,0)($\neq$ 0$\neq$ 0)( ≠ 0 , ≠ 0 ). Interestingly, there is a topological phase transition that occurs at $m_{0}=0$ for the second NN model only. The above model preserves particle-hole symmetry (PHS) $P=A\mathcal{K}$ and CS $C=\sigma_{x}s_{y}$ such that $PH_{0}(\boldsymbol{\mathrm{k}})P^{-1}=-H_{0}(-\boldsymbol{\mathrm{k}})$ and $CH_{0}(\boldsymbol{\mathrm{k}})C^{-1}=-H_{0}(\boldsymbol{\mathrm{k}})$ with $A=\sigma_{x}s_{z}$ and $\mathcal{K}$ being the complex-conjugation. Importantly, the time-reversal symmetry $T=B\mathcal{K}$ is broken when $\lambda^{f}_{1,2}\neq 0$ such that $TH_{0}(\boldsymbol{\mathrm{k}})T^{-1}\neq H_{0}(-\boldsymbol{\mathrm{k}})$ with $B=\sigma_{0}s_{y}$.
@@
Having demonstrated the physics of Hermitian SOTI model, we now focus on the NH version of the above. We associate the NH effect to the spin-orbit coupling part of the above Hamiltonian: $H_{\gamma}(\boldsymbol{\mathrm{k}})=H_{0}(\boldsymbol{\mathrm{k}})+i\gamma(\Gamma_{1}+\Gamma_{2})$ resulting in $H^{\dagger}_{\gamma}(\boldsymbol{\mathrm{k}})\neq H_{\gamma}(\boldsymbol{\mathrm{k}})$. The non-Hermiticity considered here can be thought of as an imaginary fictitious Zeeman field Yao et al. (2018); Ghosh and Nag (2022). Notice that $H_{\gamma}(\boldsymbol{\mathrm{k}})$ preserves PHS, generated by $A$ as follows $AH^{*}_{\gamma}(\boldsymbol{\mathrm{k}})A^{-1}=-H_{\gamma}(-\boldsymbol{\mathrm{k}})$. The CS continues to be preserved as $CH_{\gamma}(\boldsymbol{\mathrm{k}})C^{-1}=-H_{\gamma}(\boldsymbol{\mathrm{k}})$. We exploit the CS to define the NH analog of quadrupole moment in real space (see latter text for discussion). In the rest of the paper, we consider $\lambda^{h}_{1,2}=1$ for the sake of simplicity.
@@
Figure 1: The energy spectra under PBC and OBC are illustrated as function of the staggered mass term $m_{0}$ in upper and lower panels, respectively, for first NN hopping. The above panels correspond to (a,d) Re[$E$], (b,e) Im[$E$] and (c,f) $|E|$. The model parameters are chosen as $\lambda^{s}_{1}=\lambda^{h}_{1}=\lambda^{f}_{1}=1.0$, and $\gamma=0.2$. Here, green lines correspond to the exceptional points obtained from the PBC case $m_{0}=\pm 2\pm\tilde{\gamma}$. The red lines, representing the exceptional phase boundary under OBC, are given by $m_{0}=\pm(2+\gamma^{2}_{1})$.

### v2

# Corner modes in non-Hermitian next-nearest-neighbor hopping model

@@ Abstract
We consider a non-Hermitian (NH) analog of a second-order topological insulator, protected by chiral symmetry, in the presence of next-nearest neighbor hopping elements to theoretically investigate the interplay beyond the first nearest neighbor hopping amplitudes and topological order away from Hermiticity. In addition to the four zero-energy corner modes present in the first nearest neighbor hopping model, we uncover that the second nearest neighbor hopping introduces another topological phase with sixteen zero-energy corner modes. Importantly, the NH effects are manifested in altering the Hermitian phase boundaries for both the models. While comparing the complex energy spectrum under open boundary conditions, and bi-orthogonalized quadrupolar winding number in real space, we resolve the apparent anomaly in the bulk boundary correspondence of the NH system as compared to the Hermitian counterpart by incorporating the effect of non-Bloch form of momentum into the mass term. The above invariant is also capable of capturing the phase boundaries between the two different topological phases where the degeneracy of the corner modes is evident, as exclusively observed for the second nearest neighbor model.
@@ I Introduction
The systems with topological band properties are identified with gapless boundary modes that are characterized by symmetry-protected topological invariants. This is known as bulk boundary correspondence (BBC) Hasan and Kane (2010); Sato and Ando (2017). The conventional BBC, being an integral part of the first-order ($n=1$) topological phase Hasan and Kane (2010); Haldane (1988); Kane and Mele (2005); Bernevig et al. (2006), is generalized for higher-order ($n>1$) topological phases in $d\geq 2$ dimensions where there exist $n_{c}=(d-n)$-dimensional boundary modes Benalcazar et al. (2017a, b); Song et al. (2017); Langbehn et al. (2017); Schindler et al. (2018); Franca et al. (2018); Wang et al. (2019); Călugăru et al. (2019); Szumniak et al. (2020); Ni et al. (2020); Xie et al. (2021); Saha et al. (2022); Zhu (2018); Liu et al. (2018); Yan et al. (2018); Wang et al. (2018a); Zhang et al. (2019a, b); Volpez et al. (2019); Yan (2019a); Ghorashi et al. (2019); Wu et al. (2020); Laubscher et al. (2020); Roy (2020); Zhang et al. (2020); Kheirkhah et al. (2021); Yan (2019b); Ahn and Yang (2020); Luo et al. (2021); Wang et al. (2018b); Ghosh et al. (2021); Roy and Juričić (2021); Li et al. (2021); Ghosh et al. (2022a, b). For example, the second-order topological insulator (SOTI) in two dimensions hosts zero-dimensional (0D) localized corner modes at zero-energy while this bulk phase is characterized by nested polarization or quadrupolar moment. Very recently, it has been reported that the number of boundary modes in a topological phase can be tuned by considering next-nearest-neighbor hopping terms DeGottardi et al. (2013); Niu et al. (2012); Xie et al. (2019); Dias and Marques (2022); Mondal and Nag as well as by implementing periodic Floquet drive Thakurathi et al. (2013); Ghosh et al. (2022a). Once the extended model continues to preserve the chiral symmetry (CS), one can characterize the new topological phase by winding numbers in odd spatial dimension  Schnyder et al. (2008); Kitaev (2009); Ryu et al. (2010). The number of degenerate zero-energy states at each boundary increases according to the enhancement of the range of the hopping amplitudes as indicated by the winding number ensuring the BBC Ryu and Hatsugai (2002); Teo and Kane (2010). It is noteworthy that the one-dimensional winding number for the first-order topological systems becomes passive in the case of even-dimensional generalizations Nakahara (2018). In contrast, the higher-order topological (HOT) phase in even spatial dimension can be characterized by an appropriately defined winding number preserving CS as a constraint Benalcazar and Cerjan (2022).
@@
In recent years, thanks to the practical realization of higher-order topological phases in meta-materials Parappurath et al. (2020); Yang et al. (2019); Malzard et al. (2015); Regensburger et al. (2012) where energy conservation no longer holds El-Ganainy et al. (2018); Denner et al. (2021) and the domain of topological quantum matter can be extended to the non-Hermitian (NH) systems. The coupling to the environment  Bergholtz and Budich (2019); Yang et al. (2021); San-Jose et al. (2016), disorder/interaction-mediated quasiparticles with finite lifetime Kozii and Fu ; Yoshida et al. (2018); Shen and Fu (2018) can effectively induce complex self energy that is modeled by an NH effective Hamiltonian  Musslimani et al. (2008); Makris et al. (2008); Yao et al. (2018); Kawabata et al. (2019); Bergholtz et al. (2021); Sone et al. (2020); Denner et al. (2021); Ghosh and Black-Schaffer (2024). Interestingly, the non-Bloch nature of the wavefunction for the NH systems renormalizes the topological mass term, thus enriching the BBC such that topological phase transitions perceived with open-boundary conditions can be explained by an appropriate bulk invariant Yao and Wang (2018); Kunst et al. (2018); Liu et al. (2019); Helbig et al. (2020); Borgnia et al. (2020); Koch and Budich (2020); Zirnstein et al. (2021); Takane (2021). The NH topological systems showcase various intriguing features such as the skin effect where the bulk states accumulate at the boundary  Yao and Wang (2018); Kunst et al. (2018); Helbig et al. (2020); Kawabata et al. (2020), exceptional points where eigenstates, corresponding to the degenerate bands, coalesce Bender (2007); Heiss (2012).
@@
Going beyond the scope of the first nearest neighbor (NN) hopping, the second NN or the next-nearest-neighbor hopping elements in Hermitian systems is found to mediate versatile topological phases where the number of zero-energy modes increases Pérez-González et al. (2019); Dias and Marques (2022); An et al. (2018); Hsu and Chen (2020). In this context, the interplay between the next-NN hopping elements and non-Hermiticity is still in its infancy as far as HOT systems are concerned Liu et al. (2019); Ghosh and Nag (2022); Liu et al. (2021, 2023); Rangi et al. (2024); Ji et al. (2024). Therefore, considering a two-dimensional (2D) NH SOTI with next-NN hopping, we, therefore, examine whether non-Hermiticity induces exceptional SOTI phases, otherwise absent in the Hermitian case, and address the following interesting questions that have not been explored so far in the literature, to the best of our knowledge: How does the BBC change in the above NH phases? Can we characterize these emerging exceptional topological phases by bi-orthogonalized non-Bloch winding number?
@@
We consider a CS-preserved generic model, hosting SOTI phases in the presence of non-Hermiticity and next-NN hopping terms. The first [second] NN Hermitian model Hamiltonian can host four [four and sixteen] zero-energy corner modes while exceptional points, caused by the NH effects, reshape topological phase boundaries as compared to the Hermitian case. As a result, we find that BBC is not only different from their Hermitian counterparts but also non-trivially modified once second NN hopping terms are included. The phase boundaries between four [one] and one [none] zero-energy modes per corner in the second [first] NN model are revised following the dressed mass term due to the non-Bloch nature of wavefunction for NH Hamiltonian. We compute the quadrupolar winding number (QWN) appropriately in real space by exploiting the CS as well as implementing the bi-orthogonalization and non-Bloch nature to lay out the phase diagram, which is in accordance with the complex spectrum under open boundary conditions (OBC).
@@
The remainder of the article is organized in the following way. We discuss the details of the tight-binding Hamiltonian in Sec. II, where both the NH first and second hopping models are demonstrated. Sec. III is devoted to the main results of this article. In particular, we discuss the result associated with the first and the second NN hopping models in Secs. III.1 and III.2, respectively. Next, we illustrate the exceptional phase diagram in Sec. III.3 by examining the QWN. We finally summarize and conclude in Sec. IV. In Appendices A and B, we discuss the spatial symmetries of our model, the effect of asymmetric hoppings, and the finite size effect with the NH second NN hopping model.
@@ II Model Hamiltonian
where, $\Gamma_{1}=\sigma_{x}s_{z}$, $\Gamma_{2}=\sigma_{y}s_{0}$, $\Gamma_{3}=\sigma_{z}s_{0}$, $\Gamma_{4}=\sigma_{x}s_{x}$. We consider the strengths of first (second) NN hopping, spin-orbit coupling, and $C_{4}$ symmetry breaking mass terms as $\lambda^{h}_{1}$, $\lambda^{s}_{1}$, and $\lambda^{f}_{1}$, ($\lambda^{h}_{2}$, $\lambda^{s}_{2}$, and $\lambda^{f}_{2}$), respectively. Here, $m_{0}$ is the staggered mass term. In what follows, we refer to the case of $\lambda_{1}\neq 0$ and $\lambda_{2}=0$ ($\lambda_{1,2}\neq 0$) as first (second) NN model. In the absence of the second NN terms i.e.,   $\lambda^{h}_{2},\lambda^{s}_{2},\lambda^{f}_{2}=0$ and $\lambda^{h}_{1},\lambda^{s}_{1},\lambda^{f}_{1}\neq 0$, the Hamiltonian $H_{0}(\boldsymbol{\mathrm{k}})$ exhibits bulk gap closing at $m_{0}=\pm 2\lambda^{h}_{1}$. One can show that the Hamiltonian $H_{0}(\boldsymbol{\mathrm{k}})$ exhibits four zero-energy corner modes arising within the regime $-2\lambda^{h}_{1}<m_{0}<2\lambda^{h}_{1}$, manifesting a SOTI. Note that, in the absence of Wilson-Dirac mass terms $\lambda^{f}_{1,2}=0$, one obtains gapless edge modes for $-2\lambda^{h}_{1}<m_{0}<2\lambda^{h}_{1}$ except for $m_{0}=0$. This indicates the fact that the first-order topological insulator phase completely transforms into the second-order topological insulator phase. Furthermore, the presence of second NN terms, i.e.,   $\lambda^{h}_{2},\lambda^{s}_{2},\lambda^{f}_{2}\neq 0$ and $\lambda^{h}_{1},\lambda^{s}_{1},\lambda^{f}_{1}\neq 0$ substantially modifies the phase boundaries. To be specific, the bulk gap closes at $m_{0}=-2\lambda^{h}_{2},0,2\lambda^{h}_{1}+2\lambda^{h}_{2}$. In this case, the system harbors four zero-energy corner modes for $-2\lambda^{h}_{2}<m_{0}<0$ while sixteen zero-energy corner modes for $0<m_{0}<2\lambda^{h}_{1}+2\lambda^{h}_{2}$. Similar to the earlier case, in the absence of Wilson-Dirac mass terms $\lambda^{f}_{1,2}=0$, one obtains the gapless edge modes for $-2\lambda^{h}_{2}<m_{0}<2\lambda^{h}_{1}+2\lambda^{h}_{2}$ except for $m_{0}=0$. Note that, this model also hosts first order topological insulator phase for $-2\lambda^{h}_{1}<m_{0}<2\lambda^{h}_{1}$ and $-2\lambda^{h}_{2}<m_{0}<2\lambda^{h}_{1}+2\lambda^{h}_{2}$, respectively, when $(\lambda^{f}_{1},\lambda^{f}_{2})=(\neq 0,0)$ and (0,0)($\neq$ 0$\neq$ 0)( ≠ 0 , ≠ 0 ). The above model preserves CS $C=\sigma_{x}s_{y}$ such that $CH_{0}(\boldsymbol{\mathrm{k}})C^{-1}=-H_{0}(\boldsymbol{\mathrm{k}})$. Importantly, the time-reversal symmetry $T=i\sigma_{0}s_{y}\mathcal{K}$ is broken when $\lambda^{f}_{1,2}\neq 0$ such that $TH_{0}(\boldsymbol{\mathrm{k}})T^{-1}\neq H_{0}(-\boldsymbol{\mathrm{k}})$ with $\mathcal{K}$ being the complex-conjugation. Note that $T^{2}=-1$ leads to the AII class with a $\mathbb{Z}_{2}$ classification of the first-order topological phase.
@@
Having demonstrated the physics of the Hermitian SOTI model, we now focus on the NH version of the above. We associate the NH effect to the spin-orbit coupling part of the above Hamiltonian: $H_{\gamma}(\boldsymbol{\mathrm{k}})=H_{0}(\boldsymbol{\mathrm{k}})+i\left(\gamma_{x}\Gamma_{1}+\gamma_{y}\Gamma_{2}\right)$ resulting in $H^{\dagger}_{\gamma}(\boldsymbol{\mathrm{k}})\neq H_{\gamma}(\boldsymbol{\mathrm{k}})$. We consider $\gamma_{x}=\gamma_{y}=\gamma$ unless mentioned otherwise. The non-Hermiticity considered here can be thought of as an imaginary fictitious Zeeman field Yao et al. (2018); Ghosh and Nag (2022). The CS continues to be preserved as $CH_{\gamma}(\boldsymbol{\mathrm{k}})C^{-1}=-H_{\gamma}(\boldsymbol{\mathrm{k}})$. We exploit the CS to define the NH analog of quadrupole moment in real space (see latter text for discussion). Note that, for a NH Hamiltonian, the CS is often also referred to as sublattice symmetry Kawabata et al. (2019). In the rest of the paper, we consider $\lambda^{h}_{1,2}=1$ for the sake of simplicity.
@@
Figure 1: The energy spectra under PBC and OBC are illustrated as a function of the staggered mass term $m_{0}$ in upper and lower panels, respectively, for first NN hopping. The above panels correspond to (a,d) Re[$E$], (b,e) Im[$E$] and (c,f) $|E|$. The model parameters are chosen as $\lambda^{s}_{1}=\lambda^{h}_{1}=\lambda^{f}_{1}=1.0$, and $\gamma=0.2$. Here, green lines correspond to the exceptional points obtained from the PBC case $m_{0}=\pm 2\pm\tilde{\gamma}$. The red lines, representing the exceptional phase boundary under OBC, are given by $m_{0}=\pm(2+\gamma^{2}_{1})$.